OpenStudy (anonymous):

please help!! asap !! fan and medal !! 1.)-(10)^-1 -1/10 -1/-1^10 1/10 10 2.)1/c^-5 c^5 5c -c^5 -5/c 3.)what is the value of y^-5/x^-3 for x=2 and y=-4? 10/3 128 -1/128 -128

OpenStudy (wcrmelissa2001):

the numbers written below each questions, are they options for the answers like A B C D?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have a couple questions but im only asking 3 at a time

OpenStudy (anonymous):

ill go ahead and type the rest..

OpenStudy (wcrmelissa2001):

ok for question 1. do you know that 10^-n = 1/10^n

OpenStudy (anonymous):

im confused on all of it thats why im asking for help...

OpenStudy (anonymous):

isnt 1 a ?

OpenStudy (wcrmelissa2001):

yeah

OpenStudy (wcrmelissa2001):

wait have you done the questions already or do you need me to explaiN?

OpenStudy (anonymous):

i havnt done them

OpenStudy (wcrmelissa2001):

oh ok but for Q1 you are right

OpenStudy (wcrmelissa2001):

For question 2 it is 1/c^-5. To get rid of this negative sign , as in make this power '-5' positive, we need to bring it to the numerator. (note that if this negative power were the numerator you would bring it to the denominator). So 1/c^-5=1*c^5/1=c^5 Get it?

OpenStudy (wcrmelissa2001):

Anyways for the last question again, like the second question, you want to make the powers positive. So you bring them to the numerator or denominator. And you would get x^3/y^5. Subbstitute in the values of x and y and you should get your answer :D